OpenStudy (anonymous):

Binomial Theorem Question - The first three therms in the binomial expansion of (a+b)^n, in ascending powers of b, are denoted by p, q and r respectively. Show that (q^2/pr)=(2n/2n-1).

OpenStudy (anonymous):

I think the first thing to do would be to how the first 3 terms \[=a ^{n}+\left(\begin{matrix}n \\ 1\end{matrix}\right)a ^{n-1}b+\left(\begin{matrix}n \\ 2\end{matrix}\right)a ^{n-2}b ^{2}\]

OpenStudy (anonymous):

Yeah, that's right..I've written out this step already but am stumped after this...

OpenStudy (anonymous):

so now we have the terms we should replace them so that p=a^n q=nC1*a^(n-1)b r=nC2*a^(n-2)b^2

OpenStudy (anonymous):

we can then put them into q^2/pr [nC1*a^(n-1)b]^2/a^n*nC2*a^(n-2)b^2

OpenStudy (anonymous):

since nC1=n we can expand it to be n^2*a^(2n-2)b^2/a^n*nC2*a(n-2)b^2 n^2*a^(2n-2)b^2/nC2*a(2n-2)b^2 n^2/nC2

OpenStudy (anonymous):

I don't understand why the a^n in the first line isn't present in the second step..

OpenStudy (anonymous):

Oh ok, now I get it..it's multiplied with a^(n-2)

OpenStudy (anonymous):

Thanks a LOT!